Tori Company uses the FIFO method in its process costing system. The first processing department, the Soldering Department, started the month with 24,000 units in its beginning work-in-process inventory that were 70% complete with respect to conversion costs. The conversion cost in this beginning work-in-process inventory was $108,150. An additional 75,700 units were started into production during the month. There were 30,000 units in the ending work-in-process inventory of the Soldering Department that were 80% complete with respect to conversion costs. A total of $572,125 in conversion costs were incurred in the department during the month.
What would be the cost per equivalent unit for conversion costs? (Round off to three decimal places.)
Answer #1

Unit completed and transferred out = 24000+75700-30000 = 69700 Units

Equivalent unit of conversion under FIFO = (24000*30%+45700*100%+30000*80%) = 76900 Units

Cost per equivalent unit of conversion = 572125/76900 = 7.440 per unit